# Kondo Screening and Indirect Magnetic Exchange through a Conventional Superconductor Studied by the Density-Matrix Renormalization Group

### v1

@@ IV.2.1 RKKY-triplet phase
For weak $J$ and small $\Delta$, the ground state is a degenerate total-spin triplet $S=1$ (green color). Here, the two impurity spins are weakly coupled to the chain and form a nonlocal spin triplet while the host electron system is in a BCS singlet state. This spin triplet remains unscreened for the gapped system at $\Delta>0$, opposed to the $\Delta=0$ case, where screening takes place via two channels of the electron system Nozières and Blandin (1980); Affleck et al. (1995).
@@
The triplet formation is seen in our DMRG calculations of the impurity-spin correlation $\langle{\boldsymbol{S}}_{1}{\boldsymbol{S}}_{2}\rangle$ shown in Fig. 4 for the same system. In fact, for $J\to 0$ and finite but sufficiently small $\Delta$, we get $\langle{\boldsymbol{S}}_{1}{\boldsymbol{S}}_{2}\rangle\to\frac{1}{4}$, i.e., ferromagnetic correlation. Furthermore, Fig. 5 shows that the local Kondo correlations $\langle{\boldsymbol{S}}_{m}{\boldsymbol{s}}_{i_{m}}\rangle$ vanish in this regime.
@@ IV.2.2 RKKY-singlet phase
Both, the ground states of the RKKY triplet and the RKKY-singlet phase are sketched in Fig. 3. The transition can be understood as a discontinuous level crossing of subgap bound states in the spirit of Refs. Žitko et al. (2010, 2011); Yao et al. (2014). We will demonstrate below (Sec. V) that, at weak $J$, the presence of the RKKY-singlet phase as well as the transition from the triplet to the singlet phase at $\Delta_{\rm c}$ can also be understood perturbatively.
The RKKY-singlet phase extends to arbitrarily large $\Delta$ at weak $J$, and with increasing $J$ is bounded from below by larger and larger pairing strengths $\Delta$. At the corresponding phase boundary $\Delta_{\rm c}(J)$, the impurity-spin correlation $\langle{\boldsymbol{S}}_{1}{\boldsymbol{S}}_{2}\rangle$ jumps from strongly antiferromagnetic (Fig. 4, blue) to slightly ferromagnetic (light red) and, for $J\to\infty$, to almost vanishing values. For strong $J$, the transition line $\Delta_{\rm c}(J)$ in Fig. 4, actually reflects the (local) competition between Kondo-singlet and BCS-singlet formation, individually for each impurity. In fact, $\Delta_{\rm c}(J)$ nestles up against the transition line $\Delta^{(1)}_{\rm c}(J)$ of the single-impurity ($M=1$) model with increasing $J$. In the single-impurity model this line separates between the “Kondo”-screened ($S=0$) and a state with unscreened ($S=\frac{1}{2}$) impurity spin, and between odd and even fermion parity $\Pi$. The single-impurity phase diagram is shown in the Appendix A, see Fig. 13.
@@ IV.2.3 Local Kondo-singlet phase
For fixed $\Delta>0$ and in the limit $J\to\infty$, one expects that both impurity spins are individually screened, i.e., that two completely local Kondo singlets (LKS) are formed at $i_{1}$ and $i_{2}$ with strongly antiferromagnetic and local Kondo correlations $\langle{\boldsymbol{S}}_{m}{\boldsymbol{s}}_{i_{m}}\rangle\to-\frac{3}{4}$ (for $m=1,2$). In fact, this can be seen in Fig. 5 (dark blue). Consequently, the impurity-spin correlation vanishes $\langle{\boldsymbol{S}}_{1}{\boldsymbol{S}}_{2}\rangle\to 0$ as $J\to\infty$ and for arbitrary values of $\Delta>0$ (Fig. 4, white).
@@
As can be seen in Figs. 4 and 5, the singlet LKS phase (white and dark blue, respectively) is found at finite $J$ for pairing strengths $\Delta$ bounded from above by a second transition line $\Delta^{\prime}_{\rm c}(J)$ (with $\Delta^{\prime}_{\rm c}(J)<\Delta_{\rm c}(J)$). It is bounded from below by a third line $\Delta_{\rm c}^{\prime\prime}(J)$. With increasing $J$ the range of pairing strengths, $\Delta^{\prime}_{\rm c}(J)<\Delta<\Delta^{\prime\prime}_{\rm c}(J)$, in which the LKS phase is realized, grows, i.e., $\Delta^{\prime}_{\rm c}(J)$ increases and $\Delta^{\prime\prime}_{\rm c}(J)$ decreases with $J$, eventually extending over the entire $\Delta$ axis in the phase diagram.
@@ IV.2.4 Partial Kondo-screened phase
where the BCS singlet states, labelled by $\overline{m}$, extend over all sites except for the site $i_{m}$, and where $|M=\pm\frac{1}{2}\rangle_{m}$ are the eigenstates of $S_{mz}$. The degenerate ground-state spin doublet $\left|\Psi\right\rangle_{+}$ has odd spatial parity $p=-$, and the states $\left|\Psi\right\rangle_{-}$ forming the antibonding doublet have even parity $p=+$. Eq. (7) immediately implies that $\langle{\boldsymbol{S}}_{1}{\boldsymbol{S}}_{2}\rangle=0$, consistent with the DMRG result in the deep PKS limit, see Fig. 4. Also the local Kondo correlation in the ground state is easily computed from Eq. (7) and turns out as $\langle{\boldsymbol{S}}_{m}{\boldsymbol{s}}_{m}\rangle=-\frac{3}{8}$ for $m=1,2$. Again, this is in perfect agreement with the numerical result, see Fig. 5.
For weak $J$ but still in the PKS phase with $\Delta<\Delta_{\rm c}(J)$, ferromagnetic impurity-spin correlations build up, as is seen in Fig. 4 around $\Delta\approx 0.2$. Across the transition from the PKS to the ferromagnetic triplet phase for weak $J$, $\langle{\boldsymbol{S}}_{1}{\boldsymbol{S}}_{2}\rangle$ evolves smoothly. Note that we continue to refer to the phase under consideration as “PKS” for simplicity, even though Eq. (7) no longer holds.
@@
For small $\Delta>0$ and weak $J$ within the PKS phase, on the other hand, the situation is less clear. The data could be consistent with a critical pairing strength $\Delta_{\rm c}(J)\to 0$ for $J\to 0$ (see Fig. 2), such that the PKS state with $\Delta<\Delta_{\rm c}(J)$ terminates exactly at the point $\Delta=0$ and $J=0$ in the phase diagram with correlations $\langle{\boldsymbol{S}}_{1}{\boldsymbol{S}}_{2}\rangle\to\frac{1}{4}$ (cf. Fig. 4) and $\langle{\boldsymbol{S}}_{m}{\boldsymbol{s}}_{m}\rangle\to 0$ (Fig. 5), thereby approaching those of the RKKY-triplet phase. A direct numerical proof, however, would require calculations for $\Delta<0.1$ and correspondingly for larger values of $L$. This regime is currently inaccessible because the ground-state entanglement increases strongly with decreasing gap size $2\Delta$. We will revisit the discussion of this parameter regime in Sec. VI.
@@ IV.3 Discussion
Partial Kondo screening, i.e., screening of only a single spin in a unit cell, has been discussed in the context of the magnetically frustrated Kondo-lattice Motome et al. (2010), for the Anderson model Hayami et al. (2011); Aulbach et al. (2015) on the triangular lattice, and the Kondo lattice on the zigzag ladder Peschke et al. (2018, 2019). As it emerges as a result of the competition between RKKY coupling and Kondo screening, the PKS phase is typically located between an (RKKY driven) magnetic phase and a paramagnetic heavy-fermion phase with strong local Kondo correlations in the respective ground-state phase diagrams.
In the context of the two-impurity-spin model, the PKS ground state was first suggested in the NRG study, Ref. Yao et al. (2014) (and called “molecular doublet”). However, there is a qualitative difference between the more generic NRG phase diagram for a flat density of states and treating $J_{\rm RKKY}$ as an independent parameter rather than an emergent coupling, opposed to the present phase diagram for the concrete one-dimensional geometry: In the phase diagram of Ref. Yao et al. (2014), the PKS phase is “detached” from the $\Delta=0$ line and extends over a limited parameter space only. Contrary, in Fig. 2, the phase appears to merge with the gapless $\Delta=0$ state in a rather wide range of exchange interactions $J$.
In the gapless model at $\Delta=0$, we do not expect a quantum-critical point at a finite $J$ but rather a smooth crossover from the RKKY to the Kondo regime with increasing $J$, see Refs. Jones et al. (1988); Jones and Varma (1989); Sakai and Shimizu (1992); Fye (1994); Affleck et al. (1995); Silva et al. (1996) and the discussion in Ref. Eickhoff et al. (2018).
@@
In fact, the topology of the phase diagrams for odd $d$ and for even $d$ remains the same as $d$ increases. In particular, there are smooth crossovers rather than first-order transitions between the different regimes for odd $d$, while for even $d$ the phase diagrams consist of the above-discussed four different phases. The most important quantitative effect is the exponentially strong decrease of $J_{\rm RKKY}=J_{\rm RKKY}(d)$ with $d=i_{1}-i_{2}$. Hence, the single-impurity limit is approached rapidly as $d$ increases. This also crucially depends on $\Delta$, as the $\Delta$ dependence of $J_{\rm RKKY}=J_{\rm RKKY}(\Delta)$ is exponential as well (cf. Sec. V). Furthermore, from the comparison between $d=2$ and $d=4$ (not shown), we conclude that the phase space taken by the PKS phase at even $d$ shrinks to zero as $d\to\infty$.
@@
For the gapped system considered here, discontinuous transitions result from level crossings between different subgap bound states upon variation of the model parameters. However, a crossing between bound states in the same $(S,q)$ sector requires the fine-tuning of three independent parameters Von Neumann and Wigner (1929). From this we conclude that level transitions in the $J$-$\Delta$ plane are not to be expected and that the phase diagram therefore only shows smooth crossovers (cf. Fig. 1).
@@ V Perturbation theory
Here, ±$\pm\varepsilon_{n}=\pm\sqrt{4t^{2}\cos^{2}(k_{n})+\Delta^{2}}$ for $k_{n}=\frac{\pi n}{L+1}$ and $n=1,...,L$ are the one-particle energies. Both, the positive and the negative eigenenergies are twofold degenerate each, except for odd $L$ at $n=(L+1)/2$, where ±$\pm\varepsilon_{n}=\pm\Delta$. We note that the limit $\Delta\to 0$ in Eq. (9) is regular and that the resulting coupling constant $\lim_{\Delta\to 0}J_{\rm RKKY}(\Delta)$ equals the standard expression for $J_{\rm RKKY}$ that is obtained by perturbation theory in $J$, but starting from the $J=\Delta=0$ ground state as a reference, see the supplemental material of Ref. Schwabe et al. (2012).
@@
For a gapped system, an exponential decrease of $J_{\rm RKKY}(d)\propto\exp(-d/\xi)$ with a length scale $\xi\propto 1/\Delta$ for large $d$ and $\Delta$ is simply the result of the tunnelling effect Bloembergen and Rowland (1955). On the other hand, for small $d$ and $\Delta$, the contributions of high-energy excitations to $J_{\rm RKKY}$ dominate and cause oscillations as in the standard RKKY theory of gapless systems. Furthermore, and plausibly, the absolute magnitude of RKKY coupling is at a maximum for $d=0$ and ferromagnetic (both spins couple to the same site) for any $\Delta$. For $\Delta\to\infty$, we have the simple analytical result $J_{\rm RKKY}(d=0)=-J^{2}/8\Delta^{2}$. Combined with the general sum rule $\sum_{i^{\prime}=1}^{L}J_{ii^{\prime}}=0$, valid for a singlet ground state, this implies that for large $d$ the coupling must be antiferromagnetic (positive coupling) to compensate for the strongly negative local term. If $\Delta$ is increased, see blue line in Fig. 6, the tunnel regime is reached earlier, i.e., at smaller $d$. For $\Delta=0.5$ (green line), the RKKY coupling is antiferromagnetic for all distances $d$, except for $d=0$. This also holds for larger $\Delta$.

### v2

@@ IV.2.1 RKKY-triplet phase
For weak $J$ and small $\Delta$, the ground state is a degenerate total-spin triplet $S=1$ (see Fig. 2 green color). Here, the two impurity spins are weakly coupled to the chain and form a nonlocal spin triplet while the host electron system is in a BCS singlet state. This spin triplet remains unscreened for the gapped system at $\Delta>0$, opposed to the $\Delta=0$ case, where screening takes place via two channels of the electron system [77, 78].
@@
The triplet formation is seen in our DMRG calculations of the impurity-spin correlation $\langle{\boldsymbol{S}}_{1}{\boldsymbol{S}}_{2}\rangle$ shown in the upper plot of Fig. 4 for the same system. In fact, for $J\to 0$ and finite but sufficiently small $\Delta$, we get $\langle{\boldsymbol{S}}_{1}{\boldsymbol{S}}_{2}\rangle\to\frac{1}{4}$, i.e., ferromagnetic correlation. Furthermore, the lower plot in the figure demonstrates that the local Kondo correlations $\langle{\boldsymbol{S}}_{m}{\boldsymbol{s}}_{i_{m}}\rangle$ vanish in this regime.
@@ IV.2.2 RKKY-singlet phase
Both, the ground states of the RKKY triplet and the RKKY-singlet phase are sketched in Fig. 3. The transition can be understood as a discontinuous level crossing of subgap bound states in the spirit of Refs. [68, 69, 70]. We will demonstrate below (Sec. V) that, at weak $J$, the presence of the RKKY-singlet phase as well as the transition from the triplet to the singlet phase at $\Delta_{\rm c}$ can also be understood perturbatively.
Figure 4: Impurity-spin correlation function $\langle{\boldsymbol{S}}_{1}{\boldsymbol{S}}_{2}\rangle$ (upper plot) and Kondo correlation function $\langle{\boldsymbol{S}}_{m}{\boldsymbol{s}}_{m}\rangle$ (lower plot) for $d=2$ (the same setup as for Fig. 2).
The RKKY-singlet phase extends to arbitrarily large $\Delta$ at weak $J$, and with increasing $J$ is bounded from below by larger and larger pairing strengths $\Delta$. At the corresponding phase boundary $\Delta_{\rm c}(J)$, the impurity-spin correlation $\langle{\boldsymbol{S}}_{1}{\boldsymbol{S}}_{2}\rangle$ jumps from strongly antiferromagnetic (Fig. 4, upper plot, blue) to slightly ferromagnetic (light red) and, for $J\to\infty$, to almost vanishing values. For strong $J$, the transition line $\Delta_{\rm c}(J)$ in Fig. 4, actually reflects the (local) competition between Kondo-singlet and BCS-singlet formation, individually for each impurity. In fact, $\Delta_{\rm c}(J)$ nestles up against the transition line $\Delta^{(1)}_{\rm c}(J)$ of the single-impurity ($M=1$) model with increasing $J$. In the single-impurity model this line separates between the “Kondo”-screened ($S=0$) and a state with unscreened ($S=\frac{1}{2}$) impurity spin, and between odd and even fermion parity $\Pi$. The single-impurity phase diagram is shown in the Appendix A, see Fig. 12.
@@ IV.2.3 Local Kondo-singlet phase
For fixed $\Delta>0$ and in the limit $J\to\infty$, one expects that both impurity spins are individually screened, i.e., that two completely local Kondo singlets (LKS) are formed at $i_{1}$ and $i_{2}$ with strongly antiferromagnetic and local Kondo correlations $\langle{\boldsymbol{S}}_{m}{\boldsymbol{s}}_{i_{m}}\rangle\to-\frac{3}{4}$ (for $m=1,2$). In fact, this can be seen in the lower plot of Fig. 4 (dark blue). Consequently, the impurity-spin correlation vanishes $\langle{\boldsymbol{S}}_{1}{\boldsymbol{S}}_{2}\rangle\to 0$ as $J\to\infty$ and for arbitrary values of $\Delta>0$ (upper plot of Fig. 4, white).
@@
As can be seen in the upper and lower plot of Fig. 4, the singlet LKS phase (white and dark blue, respectively) is found at finite $J$ for pairing strengths $\Delta$ bounded from above by a second transition line $\Delta^{\prime}_{\rm c}(J)$ (with $\Delta^{\prime}_{\rm c}(J)<\Delta_{\rm c}(J)$). It is bounded from below by a third line $\Delta_{\rm c}^{\prime\prime}(J)$. With increasing $J$ the range of pairing strengths, $\Delta^{\prime}_{\rm c}(J)<\Delta<\Delta^{\prime\prime}_{\rm c}(J)$, in which the LKS phase is realized, grows, i.e., $\Delta^{\prime}_{\rm c}(J)$ increases and $\Delta^{\prime\prime}_{\rm c}(J)$ decreases with $J$, eventually extending over the entire $\Delta$ axis in the phase diagram.
@@ IV.2.4 Partial Kondo-screened phase
where the BCS singlet states, labelled by $\overline{m}$, extend over all sites except for the site $i_{m}$, and where $|M=\pm\frac{1}{2}\rangle_{m}$ are the eigenstates of $S_{mz}$. The degenerate ground-state spin doublet $\left|\Psi\right\rangle_{+}$ has odd spatial parity $p=-$, and the states $\left|\Psi\right\rangle_{-}$ forming the antibonding doublet have even parity $p=+$. Eq. (7) immediately implies that $\langle{\boldsymbol{S}}_{1}{\boldsymbol{S}}_{2}\rangle=0$, consistent with the DMRG result in the deep PKS limit, see Fig. 4 (upper plot). Also the local Kondo correlation in the ground state is easily computed from Eq. (7) and turns out as $\langle{\boldsymbol{S}}_{m}{\boldsymbol{s}}_{m}\rangle=-\frac{3}{8}$ for $m=1,2$. Again, this is in perfect agreement with the numerical result, see the lower plot in Fig. 4.
For weak $J$ but still in the PKS phase with $\Delta<\Delta_{\rm c}(J)$, ferromagnetic impurity-spin correlations build up, as is seen in Fig. 4 (upper plot) around $\Delta\approx 0.2$. Across the transition from the PKS to the ferromagnetic triplet phase for weak $J$, $\langle{\boldsymbol{S}}_{1}{\boldsymbol{S}}_{2}\rangle$ evolves smoothly. Note that we continue to refer to the phase under consideration as “PKS” for simplicity, even though Eq. (7) no longer holds.
@@
For small $\Delta>0$ and weak $J$ within the PKS phase, on the other hand, the situation is less clear. The data could be consistent with a critical pairing strength $\Delta_{\rm c}(J)\to 0$ for $J\to 0$ (see Fig. 2), such that the PKS state with $\Delta<\Delta_{\rm c}(J)$ terminates exactly at the point $\Delta=0$ and $J=0$ in the phase diagram with correlations $\langle{\boldsymbol{S}}_{1}{\boldsymbol{S}}_{2}\rangle\to\frac{1}{4}$ and $\langle{\boldsymbol{S}}_{m}{\boldsymbol{s}}_{m}\rangle\to 0$ (cf. Fig. 4), thereby approaching those of the RKKY-triplet phase. A direct numerical proof, however, would require calculations for $\Delta\ll 0.1$ and correspondingly for much larger values of $L$. We will revisit the discussion of this parameter regime in Sec. VI.
@@ IV.3 Discussion
Partial Kondo screening, i.e., screening of only a single spin in a unit cell, has been discussed in the context of the magnetically frustrated Kondo-lattice [79], for the Anderson model [80, 81] on the triangular lattice, and the Kondo lattice on the zigzag ladder [82, 83]. As it emerges as a result of the competition between RKKY coupling and Kondo screening, the PKS phase is typically located between an (RKKY driven) magnetic phase and a paramagnetic heavy-fermion phase with strong local Kondo correlations in the respective ground-state phase diagrams.
In the context of the two-impurity-spin model, the PKS ground state was first suggested in the NRG study, Ref. [70] (and called “molecular doublet”). However, there is a qualitative difference between the more generic NRG phase diagram for a flat density of states and treating $J_{\rm RKKY}$ as an independent parameter rather than an emergent coupling, opposed to the present phase diagram for the concrete one-dimensional geometry: In the phase diagram of Ref. [70], the PKS phase is “detached” from the $\Delta=0$ line and extends over a limited parameter space only. Contrary, in Fig. 2, the phase appears to merge with the gapless $\Delta=0$ state in a rather wide range of exchange interactions $J$.
In the gapless model at $\Delta=0$, we do not expect a quantum-critical point at a finite $J$ but rather a smooth crossover from the RKKY to the Kondo regime with increasing $J$, see Refs. [9, 84, 85, 86, 78, 87] and the discussion in Ref. [23].
@@
In fact, the topology of the phase diagrams for odd $d$ and for even $d$ remains the same as $d$ increases. In particular, there are smooth crossovers rather than first-order transitions between the different regimes for odd $d$, while for even $d$ the phase diagrams consist of the above-discussed four different phases. The most important quantitative effect is the exponentially strong decrease of $J_{\rm RKKY}=J_{\rm RKKY}(d)$ with $d=i_{1}-i_{2}$ for large $d$. Hence, the single-impurity limit is approached rapidly as $d$ increases. This also crucially depends on $\Delta$, as the $\Delta$ dependence of $J_{\rm RKKY}=J_{\rm RKKY}(\Delta)$ is exponential as well, for large $\Delta$ (cf. Sec. V). Furthermore, from the comparison between $d=2$ and $d=4$ (not shown), we conclude that the phase space taken by the PKS phase at even $d$ shrinks to zero as $d\to\infty$.
@@
For the gapped system considered here, discontinuous transitions result from level crossings between different subgap bound states upon variation of the model parameters. However, a crossing between bound states in the same $(S,q)$ sector requires the fine-tuning of three independent parameters [88]. From this we conclude that level transitions in the $J$-$\Delta$ plane are not to be expected and that the phase diagram therefore only shows smooth crossovers (cf. Fig. 1).
@@ V Perturbation theory
Here, ±$\pm\varepsilon_{n}=\pm\sqrt{4t^{2}\cos^{2}(k_{n})+\Delta^{2}}$ for $k_{n}=\frac{\pi n}{L+1}$ and $n=1,...,L$ are the one-particle energies. Both, the positive and the negative eigenenergies are twofold degenerate each, except for odd $L$ at $n=(L+1)/2$, where ±$\pm\varepsilon_{n}=\pm\Delta$. We note that the limit $\Delta\to 0$ in Eq. (9) is regular and that the resulting coupling constant $\lim_{\Delta\to 0}J_{\rm RKKY}(\Delta)$ equals the standard expression for $J_{\rm RKKY}$ that is obtained by perturbation theory in $J$, but starting from the $J=\Delta=0$ ground state as a reference, see the supplemental material of Ref. [36].
For a gapped system, an exponential decrease of $J_{\rm RKKY}(d)\propto\exp(-d/\xi)$ with a length scale $\xi\propto 1/\Delta$ for large $d$ and $\Delta$ is simply the result of the tunnelling effect [35]. On the other hand, for small $d$ and $\Delta$, the contributions of high-energy excitations to $J_{\rm RKKY}$ dominate and cause oscillations as in the standard RKKY theory of gapless systems. Furthermore, and plausibly, the absolute magnitude of RKKY coupling is at a maximum for $d=0$ and ferromagnetic (both spins couple to the same site) for any $\Delta$. For $\Delta\to\infty$, we have the simple analytical result $J_{\rm RKKY}(d=0)=-J^{2}/8\Delta^{2}$. Combined with the general sum rule $\sum_{i^{\prime}=1}^{L}J_{ii^{\prime}}=0$, valid for a singlet ground state, this implies that for large $d$ the coupling must be antiferromagnetic (positive coupling) to compensate for the strongly negative local term. If $\Delta$ is increased, see blue line in Fig. 5, the tunnel regime is reached earlier, i.e., at smaller $d$. For $\Delta=0.5$ (green line), the RKKY coupling is antiferromagnetic for all distances $d$, except for $d=0$. This also holds for larger $\Delta$.
@@
It is instructive to compare the results for quantum impurity spins $\frac{1}{2}$ with approaches, where the impurity spins are treated as classical [89, 90], and where the crossover from oscillatory distance dependence of the inter-impurity coupling, at small $d$ and small $\Delta$, to antiferromagnetic coupling and exponential decay, at large $d$ and $\Delta$, is explained perturbatively in $J$. In Ref. 91 it is furthermore demonstrated by nonperturbative means that the presence of YSR bound states [51, 52, 53] and their $J$-dependent position in the gap pushes this crossover to significantly smaller distances. Here, we find the same enhanced antiferromagnetic exchange within DMRG for the quantum-spin case. By comparison with the perturbative RKKY results and due to the almost perfect agreement with the DMRG data, we conclude that there is no particularly nonperturbative mechanism dominating over the perturbative RKKY coupling, at least for the parameter ranges considered. However, consistent with Ref. 91, there is a nonperturbative contribution to the enhancement of the antiferromagnetic exchange, i.e., $J^{\rm(DMRG)}_{\rm RKKY}/J^{2}$ is $J$ dependent and stronger for stronger $J$, see the results for $J=0.1$ and $J=1.5$ in Fig. 7 for $\Delta>0.5$.